# Frobenius intertwiners for $q$-difference equations

### v1

@@ Abstract.
In this note we consider a class of $q$-hypergeometric equations describing the quantum difference equation for $X=T^{*}\mathbb{P}^{1}$ We show that over $\mathbb{Q}_{p}$ these equations are equipped with the Frobenius action $z\to z^{p}$. We obtain an explicit formula for the constant term of the Frobenius intertwiner in therms of $p$ adic $q$ gamma function of Koblitz. In the limit $q\to 1$ our formula degenerates to the Frobenius intertwiners for the hypergeometric differential equations discovered by B.Dwork.
@@
Key words: Frobenius intertwiners, $q$ hypergeometric functions, $p$ adic $q$ Gamma function
@@
In this section we let $\mathbb{K}=\mathbb{C}$. In the next section we switch to the field of $p$-adic numbers.
1.1.
Let $D$ be the $q$-shift operator acting by $Df(z)=f(zq)$. We consider a $q$-holonomic module $\mathcal{H}({\boldsymbol{u}},\hbar,q,z)$ with $\mathbb{K}^{\times}$-valued parameters ${\boldsymbol{u}}=(u_{1},\dots,u_{n})$, $\hbar$ given by the following $q$-difference equation
(1.1) $\Big{(}\prod\limits_{i=1}^{n}\dfrac{1-u^{-1}_{i}D}{1-\hbar q^{-1}u^{-1}_{i}D}\Big{)}F(z)=zF(z).$
This $q$-difference equation can be written in a more standard form:
@@
(1.4) $F_{i}({\boldsymbol{u}},\hbar,q,z)=\sum\limits_{d=0}^{\infty}\,\Big{(}\prod\limits_{j=1}^{n}\,\dfrac{(u_{i}\hbar/u_{j})_{d}}{(u_{i}q/u_{j})_{d}}\Big{)}\,z^{d}.$
Set $e_{i}(z)=\exp\Big{(}{\frac{\ln(z)\ln(u_{i})}{\ln(q)}}\Big{)},$ so that $e_{i}(zq)=e_{i}(z)u_{i}$. One verifies that the functions
(1.5) $\tilde{F}_{i}({\boldsymbol{u}},\hbar,q,z)=e_{i}(z)F_{i}({\boldsymbol{u}},\hbar,q,z)$
satisfy (1.1) for all $i=1,\dots,n$. Thus, $\tilde{F}_{i}({\boldsymbol{u}},\hbar,q,z)$ give a basis of solutions for $\mathcal{H}({\boldsymbol{u}},\hbar,q,z)$.
1.3.
@@
(1.10) ${\Psi}_{i}({\boldsymbol{u}},\hbar,q,z)=\left(\begin{array}[]{l}\tilde{F}_{i}({\boldsymbol{u}},\hbar,q,z)\\ \tilde{F}_{i}({\boldsymbol{u}},\hbar,q,zq)\\ \dots\\ \tilde{F}_{i}({\boldsymbol{u}},\hbar,q,zq^{n-1})\end{array}\right)$
@@
(1.11) ${\Psi}_{i}({\boldsymbol{u}},\hbar,q,zq)=\mathbf{M}_{\mathcal{L}}({\boldsymbol{u}},\hbar,q,z)\,{\Psi}_{i}({\boldsymbol{u}},\hbar,q,z)$
for all $i=1,\dots,n$, where $\mathbf{M}_{\mathcal{L}}({\boldsymbol{u}},\hbar,q,z)$ is the companion matrix for the characteristic polynomial of (1.3):
$\mathbf{M}_{\mathcal{L}}({\boldsymbol{u}},\hbar,q,z)=\left(\begin{array}[]{cccccc}0&1&0&0&\dots&0\\ 0&0&1&0&\dots&0\\ 0&0&0&1&\dots&0\\ ..&..&..&..&..&..\\ \frac{\bar{\alpha}_{n}(z-1)}{1-\hbar^{n}z}&\frac{\bar{\alpha}_{n-1}(z\hbar-1)}{1-\hbar^{n}z}&\frac{\bar{\alpha}_{n-2}(z\hbar^{2}-1)}{\hbar^{n}z-1}&\frac{\bar{\alpha}_{n-3}(z\hbar^{3}-1)}{1-\hbar^{n}z}&\dots&\frac{\bar{\alpha}_{1}(z\hbar^{n-1}-1)}{1-\hbar^{n}z}\\ \end{array}\right)$
@@
${\Psi}({\boldsymbol{u}},\hbar,q,z)=\Big{(}{\Psi}_{1}({\boldsymbol{u}},\hbar,q,z),\dots,{\Psi}_{n}({\boldsymbol{u}},\hbar,q,z)\Big{)}.$
with $i$-th column given by ${\Psi}_{i}({\boldsymbol{u}},\hbar,q,z)$ is a fundamental solution matrix of the $q$-hypergeometric system $\mathcal{H}({\boldsymbol{u}},\hbar,q,z)$. Explicitly, it has the following matrix elements
(1.12) ${\Psi}_{i,j}({\boldsymbol{u}},\hbar,q,z)=\tilde{F}_{j}({\boldsymbol{u}},\hbar,q,zq^{i-1}).$
Taking a matrix with columns given by $\mathbb{K}$ linear combinations of the columns of $\tilde{\Psi}({\boldsymbol{u}},\hbar,q,z)$ produces new fundamental solution matrix. Thus other fundamental solution matrices are of the form
${\Psi}({\boldsymbol{u}},\hbar,q,z)\Lambda,\ \ \Lambda\in\mathsf{GL}_{n}(\mathsf{\mathbb{K}}).$
@@
(1.13) ${\Psi}({\boldsymbol{u}},\hbar,q,z)={\Psi}^{\prime}({\boldsymbol{u}},\hbar,q,z)\textsf{E}(z).$
@@
2. Vertex of $X=T^{*}\mathbb{P}^{n-1}$
The $q$-hypergeometric system $\mathcal{H}({\boldsymbol{u}},\hbar,q,z)$ describes equivariant quantum $K$-theory of the cotangent bundles over projectile spaces. This set up will be the most convenient to us and we briefly describe it in this section.
2.1.
Let (×$(\mathbb{C}^{\times})^{n}$ be the torus acting on $\mathbb{C}^{n}$ by scaling the coordinate lines with characters $u_{1},\dots,u_{n}$. We have an induced action of this torus on the cotangent bundle over projectile space $X=T^{*}\mathbb{P}(\mathbb{C}^{n})$. Let $T=(\mathbb{C}^{\times})^{n}\times\mathbb{C}^{\times}$ be a bigger torus acting on $X$: the first factor (×$(\mathbb{C}^{\times})^{n}$ acts as before and the second $\mathbb{C}^{\times}$ scales the cotangent directions with a character $\hbar^{-1}$. Recall that the $T$-equivariant K-theory of $X$ is a ring:
@@ 2. Vertex of $X=T^{*}\mathbb{P}^{n-1}$
where $L$ is the class of the line bundle $\mathcal{O}(1)$ and abusing notations we denote by the same symbols $u_{1},\dots,u_{n}$ the trivial equivariant line bundles associated to the characters $u_{i}$.
In this language, the additive parameters $a_{i}$, $h$ may be thought of as the first Chern classes of the line bundles $u_{i}$ and $\hbar$. In the same way, let $x$ be the first Chern class of $L$, i.e., we define $L=q^{x}$.
2.2.
@@
(2.2) $\Phi(\boldsymbol{a},h,x,q,z)=\Big{(}\dfrac{(q^{h},q)_{\infty}}{(q,q)_{\infty}}\Big{)}^{\!\!n}\,\prod\limits_{i=1}^{n}\dfrac{(q^{\,x+1-a_{i}},q)_{\infty}}{(q^{\,x+h-a_{i}},q)_{\infty}}\,z^{\,x}\in\hat{K}_{T}(X),$
where $(u,q)_{\infty}=\prod\limits_{m=0}^{\infty}\,(1-uq^{m})$ is the reciprocal of the $q$ gamma function. The normalized vertex function of $X$ is a generating functions counting equivariant quasimaps to quiver varieties, see Section 7.2 of [Oko17] for definitions. For $X$ this functions has the following form:
(2.3) $\widetilde{V}(\boldsymbol{a},x,h,q,z)=\sum\limits_{d=0}^{\infty}\,\Phi(\boldsymbol{a},h,x+d,q,z)\,\in\hat{K}_{T}(X)[[z]]$
@@
$P({\boldsymbol{u}},\hbar,z)\,\widetilde{V}(\boldsymbol{a},x,h,q,z)=(L-u_{1})(L-u_{2})\dots(L-u_{n})\,\Phi(\boldsymbol{a},h,x,q,z)$
where $P({\boldsymbol{u}},\hbar,z)$ is the $q$-difference operator given by (1.3). This means that in (2.1) the vertex function solves the $q$-hypergeometric equation (1.2).
2.3.
The components of the class (2.3) in any chosen basis of $K$-theory (2.1) provide a basis of solutions of the $q$-hypergeometric system $\mathcal{H}({\boldsymbol{u}},\hbar,q,z)$. For example, one convenient basis consists of the $K$-theory classes of the $T$-fixed points $X^{T}$. The components of the class (2.3) in this basis are obtained by the substitution $L=u_{i}$ or, in the additive notations, by the substitution $x=a_{i}$. In this way we obtain the basis of solutions given, up to a normalization multiple, by the hypergeometric series (1.5):
(2.4) $\widetilde{V}(\boldsymbol{a},a_{i},h,q,z)=\Phi(\boldsymbol{a},h,a_{i},q,z)\tilde{F}_{i}(\boldsymbol{a},h,q,z).$
Therefore, the vertex function (2.3) provides a uniform “basis free” description of solutions for the $q$-hypergeometric system $\mathcal{H}({\boldsymbol{u}},\hbar,q,z)$.
2.4.
@@
$\tilde{\Psi}(\boldsymbol{a},h,q,z)=\Big{(}\tilde{\Psi}_{1}(\boldsymbol{a},h,q,z),\dots,\tilde{\Psi}_{n}(\boldsymbol{a},h,q,z)\Big{)},$
where
$\tilde{\Psi}_{i}(\boldsymbol{a},h,q,z)={\Psi}_{i}(\boldsymbol{a},h,q,z)\Phi(\boldsymbol{a},h,a_{i},q,z)$
and ${\Psi}_{i}(\boldsymbol{a},h,q,z)$ are given by (1.10). We can also rewrite the last formula simply as
$\tilde{\Psi}(\boldsymbol{a},h,q,z)={\Psi}(\boldsymbol{a},h,q,z)\Phi(\boldsymbol{a},h,x,q,z)$
where $\Phi(\boldsymbol{a},h,x,q,z)$ is now understood as the operator of multiplication by the gamma class (2.2) in the equivariant K - theory. In particular, in the basis of the $T$-fixed points this operator is represented by the diagonal matrix with eigenvalues the $\Phi(\boldsymbol{a},h,a_{i},q,z)$, $i=1,\dots,n$.
2.5.
The normalized fundamental solution matrices $\tilde{\Psi}(\boldsymbol{a},h,q,z)$ also have an enumerative meaning: these are the matrices of the capping operator for $X$, which is the partition function of the relative quasimaps to $X$, see Section 7.4 in [Oko17] for the definitions.
An important input from the enumerative geometry we need for our goals is the existence of the shift operators. The operators describe $q$-difference equations satisfied by the capping operator in the equivariant parameters $\boldsymbol{a}$ and $h$. These equations exist in general for an arbitrary Nakajima variety. Applied to $X=T^{*}\mathbb{P}^{n-1}$ this results can be formulated as follows:
@@
The normalized fundamental solution matrices satisfy the following system of difference equations
$\tilde{\Psi}(a_{1},\dots,a_{i}+1,\dots a_{n},h,q,z)=S^{a_{i}}(\boldsymbol{a},h,q,z)\tilde{\Psi}(a_{1},\dots,a_{i},\dots a_{n},h,q,z),$
$\tilde{\Psi}(a_{1},\dots,a_{i},\dots a_{n},h+1,q,z)=S^{h}(\boldsymbol{a},h,q,z)\tilde{\Psi}(a_{1},\dots,a_{i},\dots a_{n},h,q,z).$

### v2

@@ Abstract.
We consider a class of $q$-hypergeometric equations describing the quantum difference equation for the cotangent bundles over projective spaces $X=T^{*}\mathbb{P}^{n-1}$ We show that over $\mathbb{Q}_{p}$ these equations are equipped with the Frobenius action $(q,z)\to(q^{p},z^{p})$. We obtain an explicit formula for the constant term of the Frobenius intertwiner in terms of the $p$-adic $q$-gamma function of Koblitz. In the limit $q\to 1$ we arrive at the Frobenius structures for the $p$-adic hypergeometric and Bessel differential equations studied by Dwork. In particular, we find closed formulas for $p$-adic constants appearing in works of Dwork and Sperber in terms of $p$-adic zeta functions.
@@
Key words: Frobenius structures, $q$-hypergeometric functions, $p$-adic $q$-Gamma function, Nakajima varieties, enumerative geometry
@@
1. Introduction
1.1. $q$-difference equations and enumerative geometry
The quantum K-theory of quiver varieties [Oko17] is governed by the quantum difference equations (QDE) [OS22]. It is expected that over $\mathbb{Q}_{p}$ these equations are equipped with a symmetry which acts on the arguments of the equations by $z\to z^{p}$. In the limit $q\to 1$ these equations reduce to the quantum differential equations of quiver varieties [MO19] while the symmetry $z\to z^{p}$ specializes to the Frobenius structures, which are well known in the theory of $p$-adic differential equations [Dw89, Ked21]. In this sense, the expected symmetry of QDE may be viewed as a $q$-deformation of the Frobenius structures.
One may hope that the $q$-deformed Frobenius structures arising this way may be of significance in arithmetic geometry, see [Sch, BS].
1.2. $q$-hypergeometric case
In this paper consider the $q$-deformed Frobenius structure for simplest examples of quiver varieties given by the cotangent bundles over projective spaces $X=T^{*}\mathbb{P}^{n-1}$. The cotangent bundle $X$ is equipped with a natural action of a torus $T=(\mathbb{C}^{*})^{n+1}$. We denote by $h$ and $\boldsymbol{a}=(a_{1},\dots,a_{n})$ the corresponding equivariant parameters of $T$. The parameter $h$ denotes the character of the one-dimensional representation $\mathbb{C}\omega$ spanned by the symplectic form $\omega\in H^{2}(X,\mathbb{C})$. This parameter plays a special role in the present paper.
In Section 2 we describe the QDE of $X$ which is nothing but the very classical $q$-hypergeometric equation with parameters $h$ and $\boldsymbol{a}$.
1.3. $q$-deformed Frobenius intertwiner
In the context of enumerative geometry it is standard to represent the fundamental solution matrix of QDE as a power series with coefficients in the equivariant $K$-theory:
(1.1) $\Psi(h,\boldsymbol{a},q,z)\in K_{T}(X)^{\otimes 2}[[z]]$
In Section 2 we fix such $\tilde{\Psi}(h,\boldsymbol{a},q,z)$ where $\tilde{}$ refers to a certain specific choice of the normalization, which involves $q$-gamma functions. The key property of this normalization, relevant to our work, is that $\tilde{\Psi}(h,\boldsymbol{a},q,z)$ satisfies a set of regular difference equations in the equivariant parameters $h$ and $\boldsymbol{a}$; see Theorem 3.1.
We consider the power series
(1.2) $\mathsf{U}(h,\boldsymbol{a},q,z)=\tilde{\Psi}(ph,p\boldsymbol{a},q,z)\tilde{\Psi}(h,\boldsymbol{a},q^{p},z^{p})^{-1}\in K_{T}(X)^{\otimes 2}[[z]]$
Fixing a basis in $K_{T}(X)$ we can represent $\mathsf{U}(h,\boldsymbol{a},q,z)$ by an $n\times n$-matrix with coefficients given by power series in $z$. Assume that
$|q-1|_{p}<1,\ \ \textrm{and}\ \ (h,a_{1},\dots,a_{n})\in\mathbb{Z}^{n+1}_{p}$
We show that under these assumptions (Theorem 5.2), $\mathsf{U}(h,\boldsymbol{a},q,z)\in\mathsf{GL}_{n}(\hat{E}_{p})$ where $\hat{E}_{p}$ denotes the field of $p$-adic analytic functions. This means that $\mathsf{U}(h,\boldsymbol{a},q,z)$ is a Frobenius intertwiner between the $q$-hypergeometric systems with parameters $(h,\boldsymbol{a},q^{p},z^{p})$ and $(ph,p\boldsymbol{a},q,z)$.
We observe that the Frobenius intertwiner $\mathsf{U}(h,\boldsymbol{a},q,z)$ has simple specializations at $q$ given by $p$-adic unit roots. If $q$ is a $p$-adic unit root of order $p^{s}$ then
$|q-1|_{p}=p^{-\frac{1}{p^{s-1}(p-1)}}<1$
We show that if $q$ is such a root (Theorem 5.8) then the Frobenius intertwiner $\mathsf{U}(h,\boldsymbol{a},q,z)$ specializes to a rational function of variables $z$ and $q^{ph},\ \ q^{pa_{1}},\dots,q^{pa_{n}}.$ In particular, these rational functions satisfy
$\mathsf{U}(h,\boldsymbol{a},q,z)=\mathsf{U}(h^{(s)},\boldsymbol{a}^{(s)},q,z)$
where $\boldsymbol{a}^{(s)}=(a_{1}^{(s)},\dots,a^{(s)}_{n})$ and $h^{(s)}$ are natural numbers defined by $h^{(s)}=h\pmod{p^{s-1}}$, $a^{(s)}_{i}=a_{i}\pmod{p^{s-1}}$, $i=1,\dots,n$. This property can be considered as a $q$-version of the Dwork’s congruences for the classical hypergeometric functions [Dw69] and the Hasse-Witt matrices [VZ].
1.4. The constant term of the Frobenius intertwiner
In Section 5, we give an explicit formula for the constant term of the Frobenius intertwiner (1.2), i.e., its specialization at $z=0$. First, for a vector bundle $\mathcal{V}$ over $X$ we define a $K$-theory valued characteristic class $\Gamma_{p,q}(\mathcal{V})\in K_{T}(X)$. This class is constructed from the $p$-adic $q$-gamma function $\Gamma_{p,q}$ defined by Koblitz [Ko80]. Second, we show that $\mathsf{U}(h,\boldsymbol{a},q,0)$ is an operator acting in $K_{T}(X)$ as the operator of multiplication by the $K$-theory class
(1.3) $\mathsf{U}(h,\boldsymbol{a},q,0)=\Gamma_{p,q}((q^{p}-q^{ph})T^{1/2}X^{(p)})$
where $T^{1/2}X$ denotes the polarization bundle of $X$ and $T^{1/2}X^{(p)}=\psi^{p}(T^{1/2}X)$ its twist by the $p$-th Adams operation.
For generic values of the equivariant parameters $a_{i}\neq a_{j}$, the equivariant $K$-theory $K_{T}(X)$ has a distinguished basis given by the $K$-theory classes of $T$-fixed points. In this basis the operator $\mathsf{U}(h,\boldsymbol{a},q,0)$ is diagonal, with the eigenvalues
$\mathsf{U}(h,\boldsymbol{a},q,0)_{i,i}=\prod\limits_{j=1}^{n}\,\dfrac{\Gamma_{p,q}(pa_{i}+ph-pa_{j})}{\Gamma_{p,q}(pa_{i}-pa_{j})\Gamma_{p,q}(ph)},\ \ i=1,\dots,n.$
We expect that (1.3) holds for the constant term of the Frobenius intertwiners for more general varieties.
1.5. Limits and specializations
The $q$-deformed Frobenius structure has several important specializations. First, the limit $q\to 1$ corresponds to reduction of the equivariant $K$-theory to the equivariant cohomology. In this limit the QDE specializes to the hypergeometric differential equation - the quantum differential equation of $X=T^{*}\mathbb{P}^{n-1}$. The $p$-adic $q$-gamma function $\Gamma_{p,q}$ specializes to the $p$-adic gamma function $\Gamma_{p}$ of Morita [Mo75]. Combining this together we find that $\left.\mathsf{U}(h,\boldsymbol{a},q,z)\right|_{q=1}$ is the Frobenius structure for the hypergeometric differential equation. Its constant term is the operator acting in the equivariant cohomology $H^{*}_{T}(X)$ as multiplication by the cohomology class $\Gamma_{p}(TX^{(p)})$ where $TX^{(p)}$ denotes the tangent bundle of $X$ twisted by $p$-th Adams operation. For generic values of the equivariant parameters this operator is diagonal in the basis of $T$-fixed points with eigenvalues
(1.4) $\ \ \ \ \ \ \small\mathsf{U}(h,\boldsymbol{a},1,0)_{i,i}=\prod\limits_{w\in\textrm{char}(T_{i}X)}\,\Gamma_{p}(pw)=\prod\limits_{j=1}^{n}\,\Gamma_{p}(pa_{i}+ph-pa_{j})\Gamma_{p}(pa_{j}-pa_{i})\ \ \ i=1,\dots,n.$
where the first product is over the $T$-characters appearing in the tangent space to $X$ at $i$-th $T$-fixed point. In this way we reproduce the results of Dwork [Dw69, Dw89] where the Frobenius intertwiners for the hypergeometric equations were discovered and also give them a geometric interpretation. We also refer to [Ked21] where a formula equivalent to (1.4) was obtained using different tools.
Second, it is well known that in the limit $h\to\infty$ the quantum differential equation of $X=T^{*}\mathbb{P}^{n-1}$ specializes to quantum differential equation of $\mathbb{P}^{n-1}$. In the most degenerate case $a_{1}=a_{2}=\dots=a_{n}$, this equation coincides with the generalized Bessel equation studied over $\mathbb{Q}_{p}$ by Sperber in [Sp80]. For $n=2$ this is the very classical Bessel equation investigated by Dwork [Dw74]. As an application, in the limit $h\to\infty$ we obtain an explicit description for the Frobenius structures of Bessel equations. In particular, we obtain closed formulas for certain $p$-adic constants appearing in [Dw74, Sp80] as values of $p$-adic zeta functions.
1.6. Enumerative definition of Frobenius intertwiner
In quantum cohomology over fields of positive characteristic one can define new classes of operators known as the quantum Steenrod operations [F96, W20]. These operations are defined enumeratively, namely as partition functions counting stable maps from $\mathbb{P}^{1}$ with $p$-marked points to $X$. The stable maps are assumed to be invariant under the cyclic group $\mathbb{Z}/p\mathbb{Z}$, which permutes the marked via rotation of the source $\mathbb{P}^{1}$. It was recently shown in [HL24] that for a large family of varieties, the quantum Steenrod operations coincide with the $p$-curvature of the quantum connection of these varieties.
We expect that this story extends naturally to the level of quantum $K$-theory of quiver varieties which is defined via equivariant count of quasimaps [Oko17]. In particular, the $q$-difference equations have natural analogs of $p$-curvature [KS]. In this context, the quantum Steenrod operations shall be lifted to the $K$-theoretic quantum Adams operations. Moreover, the Frobenius intertwiner $\mathsf{U}(h,\boldsymbol{a},q,z)$ itself can be understood as a partition function counting equivariant quasimaps from $\mathscr{C}=\mathbb{P}^{1}$ to $X$ with two relative boundary conditions at $0,\infty\in\mathscr{C}$, where the relative quasimap moduli space at $\infty$ is assumed to be $\mathbb{Z}/p\mathbb{Z}$-equivariant in an appropriate sense.
When $q$ is specialized at $p$-th unit roots, the partition function $\mathsf{U}(h,\boldsymbol{a},q,z)$ is an operator which conjugates the $p$-curvature of the $q$-difference connection to the Frobenius twist of this connection. In particular, these objects have the same spectrum [KS]. This hints at the arithmetic significance of $q$-deformed Frobenius intertwiners with $q$ specialized at the unit roots.
We briefly touch on these ideas in Section 9.
@@ 2. $q$-hypergeometric functions
In this section we work over $\mathbb{C}$. In the next section we switch to the field of $p$-adic numbers.
2.1.
Let $D$ be the $q$-shift operator acting by $Df(z)=f(zq)$. We consider a $q$-difference equation:
@@
(2.3) $f_{i}(\hbar,{\boldsymbol{u}},q,z)=\sum\limits_{d=0}^{\infty}\,\Big{(}\prod\limits_{j=1}^{n}\,\dfrac{(u_{i}\hbar/u_{j})_{d}}{(u_{i}q/u_{j})_{d}}\Big{)}\,z^{d},$
where $(x,q)_{d}:=(1-x)(1-xq)\dots(1-xq^{d-1})$. Set $e_{i}(z)=\exp\Big{(}{\frac{\ln(z)\ln(u_{i})}{\ln(q)}}\Big{)},$ so that $e_{i}(zq)=e_{i}(z)u_{i}$. One verifies that the functions
(2.4) ${F}_{i}(\hbar,{\boldsymbol{u}},q,z)=e_{i}(z)f_{i}(\hbar,{\boldsymbol{u}},q,z)$
satisfy (2.1) for all $i=1,\dots,n$. Thus, ${F}_{i}({\boldsymbol{u}},\hbar,q,z)$ give a basis of solutions to (2.1).
2.3.
@@
(2.9) ${\Psi}_{i}(\hbar,{\boldsymbol{u}},q,z)=\left(\begin{array}[]{l}{F}_{i}(\hbar,{\boldsymbol{u}},q,z)\\ {F}_{i}(\hbar,{\boldsymbol{u}},q,zq)\\ \dots\\ {F}_{i}(\hbar,{\boldsymbol{u}},q,zq^{n-1})\end{array}\right)$
@@
(2.10) ${\Psi}_{i}(\hbar,{\boldsymbol{u}},q,zq)=\mathbf{M}(\hbar,{\boldsymbol{u}},q,z)\,{\Psi}_{i}(\hbar,{\boldsymbol{u}},q,z)$
for all $i=1,\dots,n$, where $\mathbf{M}(\hbar,{\boldsymbol{u}},q,z)$ is the companion matrix for the characteristic polynomial of (2.2):
$\mathbf{M}(\hbar,{\boldsymbol{u}},q,z)=\left(\begin{array}[]{cccccc}0&1&0&0&\dots&0\\ 0&0&1&0&\dots&0\\ 0&0&0&1&\dots&0\\ ..&..&..&..&..&..\\ \frac{\bar{\alpha}_{n}(z-1)}{1-\hbar^{n}z}&\frac{\bar{\alpha}_{n-1}(z\hbar-1)}{1-\hbar^{n}z}&\frac{\bar{\alpha}_{n-2}(z\hbar^{2}-1)}{\hbar^{n}z-1}&\frac{\bar{\alpha}_{n-3}(z\hbar^{3}-1)}{1-\hbar^{n}z}&\dots&\frac{\bar{\alpha}_{1}(z\hbar^{n-1}-1)}{1-\hbar^{n}z}\\ \end{array}\right)$
@@
${\Psi}(\hbar,{\boldsymbol{u}},q,z)=\Big{(}{\Psi}_{1}(\hbar,{\boldsymbol{u}},q,z),\dots,{\Psi}_{n}(\hbar,{\boldsymbol{u}},q,z)\Big{)}.$
with $i$-th column given by ${\Psi}_{i}({\boldsymbol{u}},\hbar,q,z)$ is a fundamental solution matrix of the $q$-hypergeometric system (2.10). Explicitly, it has the following matrix elements
(2.11) ${\Psi}_{i,j}(\hbar,{\boldsymbol{u}},q,z)={F}_{j}(\hbar,{\boldsymbol{u}},q,zq^{i-1}).$
Any matrix with columns given by $\mathbb{C}$ linear combinations of the columns of ${\Psi}({\boldsymbol{u}},\hbar,q,z)$ produces a new fundamental solution matrix. Thus, other (analytic) fundamental solution matrices are of the form
${\Psi}(\hbar,{\boldsymbol{u}},q,z)\Lambda,\ \ \Lambda\in\mathsf{GL}_{n}(\mathsf{\mathbb{C}}).$
@@
(2.12) ${\Psi}(\hbar,{\boldsymbol{u}},q,z)={\Psi}^{\prime}(\hbar,{\boldsymbol{u}},q,z)\textsf{E}(z).$
@@
In the additive notations we write:
(2.14) $\Psi_{i,j}(h,\boldsymbol{a},q,z)=f_{i}(h,\boldsymbol{a},q,zq^{i-1})z^{a_{j}}q^{a_{j}(i-1)}$
where we denote by $\Psi_{i,j}(h,\boldsymbol{a},q,z)$ and $f_{i}(h,\boldsymbol{a},q,z)$ the functions (2.11) and (2.3) with parameters (2.13) respectively.
3. Vertex function of $X=T^{*}\mathbb{P}^{n-1}$
QDE (2.10) describes the equivariant quantum $K$-theory of the cotangent bundles over projectile spaces [OS22]. We briefly overview the connection in this section.
3.1.
Let (×$(\mathbb{C}^{\times})^{n}$ be the torus acting on $\mathbb{C}^{n}$ by scaling the coordinate lines with characters $u_{1},\dots,u_{n}$. We have an induced action of this torus on the cotangent bundle over projectile space $X=T^{*}\mathbb{P}(\mathbb{C}^{n})$. Let $T=(\mathbb{C}^{\times})^{n}\times\mathbb{C}^{\times}$ be a bigger torus acting on $X$: the first factor (×$(\mathbb{C}^{\times})^{n}$ acts as before and the second $\mathbb{C}^{\times}$ scales the cotangent directions with a character $\hbar^{-1}$. The $T$-equivariant $K$-theory of $X$ is the ring:
@@ 3. Vertex function of $X=T^{*}\mathbb{P}^{n-1}$
where $L$ is the class of the line bundle $\mathcal{O}(1)$. Abusing notations we denote by the same symbols $u_{1},\dots,u_{n}$ the trivial equivariant line bundles associated to the characters $u_{i}$. In this language, the additive parameters $a_{i}$, $h$ may be thought of as the first Chern classes of the line bundles $u_{i}$ and $\hbar$. Similarly, let $x$ be the first Chern class of $L$, then in our conventions $L=q^{x}$.
3.2.
@@
(3.2) $\Phi(h,\boldsymbol{a},x,q,z)=\Big{(}\dfrac{(q^{h},q)_{\infty}}{(q,q)_{\infty}}\Big{)}^{\!\!n}\,\prod\limits_{i=1}^{n}\dfrac{(q^{\,x+1-a_{i}},q)_{\infty}}{(q^{\,x+h-a_{i}},q)_{\infty}}\,z^{\,x}\in\hat{K}_{T}(X),$
where $(u,q)_{\infty}=\prod\limits_{m=0}^{\infty}\,(1-uq^{m})$ is the reciprocal of the $q$ gamma function. The normalized vertex function of $X$ is a generating function counting equivariant quasimaps to quiver varieties, see Section 7.2 of [Oko17] for definitions. For $X$ the vertex function has the following form:
(3.3) $\widetilde{V}(h,\boldsymbol{a},x,q,z)=\sum\limits_{d=0}^{\infty}\,\Phi(h,\boldsymbol{a},x+d,q,z)\,\in\hat{K}_{T}(X)[[z]]$
@@
$P({\boldsymbol{u}},\hbar,z)\,\widetilde{V}(h,\boldsymbol{a},x,q,z)=(L-u_{1})\dots(L-u_{n})\,\Phi(h,\boldsymbol{a},x,q,z)$
where $P({\boldsymbol{u}},\hbar,z)$ is the $q$-difference operator given by (2.2). Note that in (3.1) the vertex function solves the $q$-hypergeometric equation (2.1).
3.3.
The components of the class (3.3) in any chosen basis of $K$-theory (3.1) provide a basis of solutions of the $q$-hypergeometric system (2.1). One convenient basis consists of the $K$-theory classes of the $T$-fixed points $X^{T}$. The components of the class (3.3) in this basis are obtained by the specialization $L=u_{i}$ or, in the additive notations, by the specialization $x=a_{i}$. In this way we obtain the basis of solutions given, up to a normalization multiple, by the hypergeometric series (2.4):
(3.4) $\widetilde{V}(h,\boldsymbol{a},a_{i},q,z)=\Phi(h,\boldsymbol{a},a_{i},q,z){F}_{i}(h,\boldsymbol{a},q,z).$
In this way, the vertex function (3.3) provides a uniform “basis free” description of the solutions to (2.1).
3.4.
@@
$\tilde{\Psi}(h,\boldsymbol{a},q,z)=\Big{(}\tilde{\Psi}_{1}(h,\boldsymbol{a},q,z),\dots,\tilde{\Psi}_{n}(h,\boldsymbol{a},q,z)\Big{)},$
with $i$-th column
$\tilde{\Psi}_{i}(h,\boldsymbol{a},q,z)={\Psi}^{\prime}_{i}(h,\boldsymbol{a},q,z)\Phi(h,\boldsymbol{a},a_{i},q,z)$
where ${\Psi}^{\prime}_{i}(h,\boldsymbol{a},q,z)$ is the analytic part of the fundamental solution matrix as defined in (2.12). We can also write the last formula simply as
(3.5) $\tilde{\Psi}(h,\boldsymbol{a},q,z)={\Psi}^{\prime}(h,\boldsymbol{a},q,z)\Phi(h,\boldsymbol{a},x,q,z)$
where $\Phi(\boldsymbol{a},h,x,q,z)$ is now understood as the operator of multiplication by the gamma class (3.2) in the equivariant $K$ theory. In particular, in the basis of the $T$-fixed points this operator is represented by the diagonal matrix with eigenvalues given by $\Phi(h,\boldsymbol{a},a_{i},q,z)$, $i=1,\dots,n$.
3.5.
The normalized fundamental solution matrix $\tilde{\Psi}(\boldsymbol{a},h,q,z)$ has an enumerative meaning: it coincides with the capping operator for $X$, which is the partition function of the relative quasimaps to $X$, see Section 7.4 in [Oko17] for the definitions. In particular, normalized as in (3.5) the fundamental solution matrix satisfies the shift equations which are the difference equations in the equivariant parameters $\boldsymbol{a}$ and $h$. These equations exist in general for an arbitrary Nakajima variety. Applied to $X=T^{*}\mathbb{P}^{n-1}$ they can be formulated as follows:
@@
The normalized fundamental solution matrix (3.5) satisfies the following system of difference equations
$\tilde{\Psi}(h,a_{1},\dots,a_{i}+1,\dots a_{n},q,z)=A_{i}(\boldsymbol{a},h,q,z)\tilde{\Psi}(h,a_{1},\dots,a_{i},\dots a_{n},q,z),$
$\tilde{\Psi}(h+1,a_{1},\dots,a_{i},\dots a_{n},q,z)=H(\boldsymbol{a},h,q,z)\tilde{\Psi}(h,a_{1},\dots,a_{i},\dots a_{n},q,z).$